Find the unknown:
x + y - 2z = 7
2x - 3y + 2z = 0
x - 2y - 3z = 3

Answer by ewatrrr(24785)

 
Hi,
Find the unknown:
x + y - 2z = 7
2x - 3y + 2z = 0
x - 2y - 3z = 3
d = 23, Dx = 79, Dy = 38, Dz = -22
x = 79/23, y = 38/23, z = -22/23